OpenStudy (kainui):

Can we say that if f(x)=x^2 and g(x)=x^3 that f(infinity)=g(infinity)?

OpenStudy (kainui):

I guess, is it completely ridiculous to say that everything converges at infinity?

OpenStudy (anonymous):

Haha interesting

OpenStudy (kainui):

Maybe, is this a true thing to say? \[\lim_{x \rightarrow \infty}\frac{ 1 }{ x^2 }=\lim_{x \rightarrow \infty}\frac{ 1 }{ x^3 }\]

OpenStudy (kainui):

See, I have this suspicion that if you were to graph them from 0 to infinity you'd get something like this: